Greeting cards are sold in packs of 6,8 and 24 cards . The minimum number of packs needed to buy exactly 110 cards is what?
Answer by Edwin McCravy(20054)


To buy the least number of packs, you'd want to buy as
many 24-packs as possible.

Since 110÷24 = 4.58333..., you would buy the whole number
part of that, which is 4 of the 24-packs. 

That would give you 24×4 or 96 greeting cards.

You would need 110-96 or 14 more greeting cards.

So you would need only 1 6-pack and 1 8-pack, since 6+8=14.

So you would buy 4 24-packs, 1 6-pack and 1 8-pack.

That's 4+1+1 = 6 packs minimum  <--answer

Edwin